OpenStudy (anonymous):

write the expression in radical notation. 27^-2/3,evaluate the expression afterwards.

OpenStudy (anonymous):

can someone help me please

OpenStudy (anonymous):

Please use brackets, you should have written 27^(-2/3). They're not the same thing! First thing is the negative exponent means you do 1 over the positive exponent. so 27^(-2/3)=1/(27^(2/3)) Next, you can change a number that is to the power of a fraction by making the numerator the exponent and the denominator the nth root. So in this case you're going to have the 3rd (denominator) root of 27 to the power of 2 (numerator). So 1/(27^(2/3))=\[\frac{ 1 }{ \sqrt[3]{27}^{2}}\] The cube root of 27 is 3, and 3^2 is 9, so your answer is 1/9.

OpenStudy (anonymous):

thanks